The observed total loss per cycle per meter plotted as a function of the peak amplitude of the magnetic field. Also shown are the calculated values of the losses assuming that all the filaments are fully coupled (solid line) and fully uncoupled (dashed line). It is well known that multifilamentary tapes can be used to reduce their a.c. losses when their filaments are uncoupled, so as to behave as many small superconductors rather than one monoblock. However, given a high enough rate of change of the magnetic field, currents can be driven across the normal matrix of a multifilament superconductor resulting in the coupling of the filaments. These currents have an associated loss which must be added to the losses of the individual filaments. The outer layer of filaments will become saturated and the next layer of filaments will begin to conduct. When the entire composite is saturated, it behaves as a single monoblock, and in this region the losses can be estimated assuming the composite is a solid superconductor. As the complete field range used in this measurement is well above this value the tape should be fully coupled. The a.c. losses in a parallel field for the limiting cases of fully uncoupled and fully coupled filaments can be calculated using Bean’s critical state model for a one dimensional slab. The penetration field is dependent upon the dimensions of the superconductor and therefore whether the filaments are fully coupled or uncoupled. For a fully uncoupled tape. The good correspondence between the measured data and the calculations for fully coupled filaments show that the filaments in the tape are fully coupled together. The measured loss per cycle per unit length of Sample B as a function of the peak amplitude of the parallel a.c. magnetic field. Also shown are the losses calculated for fully coupled filaments (dashed line), fully uncoupled filaments (dotted line) and fully uncoupled filaments plus the coupling current losses (solid line). It shows the observed loss per cycle per unit length plotted (symbols) as a function of the peak magnetic field. It can be seen that these calculated values do not correspond to the measured values. Calculations were also undertaken assuming that the filaments had completely decoupled to give 75 individual conductors with cross-section of 210×20 m. It is seen that these calculations underestimate the actual measured values. It is known however, that the coupling currents which flow through the silver matrix also produce losses and therefore may be responsible for this discrepancy.

The observed total loss per cycle per meter plotted as a function of the peak amplitude of the magnetic field. Also shown are the calculated values of the losses assuming that all the filaments are fully coupled (solid line) and fully uncoupled (dashed line). It is well known that multifilamentary tapes can be used to reduce their a.c. losses when their filaments are uncoupled, so as to behave as many small superconductors rather than one monoblock. However, given a high enough rate of change of the magnetic field, currents can be driven across the normal matrix of a multifilament superconductor resulting in the coupling of the filaments. These currents have an associated loss which must be added to the losses of the individual filaments. The outer layer of filaments will become saturated and the next layer of filaments will begin to conduct. When the entire composite is saturated, it behaves as a single monoblock, and in this region the losses can be estimated assuming the composite is a solid superconductor. As the complete field range used in this measurement is well above this value the tape should be fully coupled. The a.c. losses in a parallel field for the limiting cases of fully uncoupled and fully coupled filaments can be calculated using Bean’s critical state model for a one dimensional slab. The penetration field is dependent upon the dimensions of the superconductor and therefore whether the filaments are fully coupled or uncoupled. For a fully uncoupled tape. The good correspondence between the measured data and the calculations for fully coupled filaments show that the filaments in the tape are fully coupled together. The measured loss per cycle per unit length of Sample B as a function of the peak amplitude of the parallel a.c. magnetic field. Also shown are the losses calculated for fully coupled filaments (dashed line), fully uncoupled filaments (dotted line) and fully uncoupled filaments plus the coupling current losses (solid line). It shows the observed loss per cycle per unit length plotted (symbols) as a function of the peak magnetic field. It can be seen that these calculated values do not correspond to the measured values. Calculations were also undertaken assuming that the filaments had completely decoupled to give 75 individual conductors with cross-section of 210×20 m. It is seen that these calculations underestimate the actual measured values. It is known however, that the coupling currents which flow through the silver matrix also produce losses and therefore may be responsible for this discrepancy.

The observed total loss per cycle per meter plotted as a function of the peak amplitude of the magnetic field. Also shown are the calculated values of the losses assuming that all the filaments are fully coupled (solid line) and fully uncoupled (dashed line). It is well known that multifilamentary tapes can be used to reduce their a.c. losses when their filaments are uncoupled, so as to behave as many small superconductors rather than one monoblock. However, given a high enough rate of change of the magnetic field, currents can be driven across the normal matrix of a multifilament superconductor resulting in the coupling of the filaments. These currents have an associated loss which must be added to the losses of the individual filaments. The outer layer of filaments will become saturated and the next layer of filaments will begin to conduct. When the entire composite is saturated, it behaves as a single monoblock, and in this region the losses can be estimated assuming the composite is a solid superconductor. As the complete field range used in this measurement is well above this value the tape should be fully coupled. The a.c. losses in a parallel field for the limiting cases of fully uncoupled and fully coupled filaments can be calculated using Bean’s critical state model for a one dimensional slab. The penetration field is dependent upon the dimensions of the superconductor and therefore whether the filaments are fully coupled or uncoupled. For a fully uncoupled tape. The good correspondence between the measured data and the calculations for fully coupled filaments show that the filaments in the tape are fully coupled together. The measured loss per cycle per unit length of Sample B as a function of the peak amplitude of the parallel a.c. magnetic field. Also shown are the losses calculated for fully coupled filaments (dashed line), fully uncoupled filaments (dotted line) and fully uncoupled filaments plus the coupling current losses (solid line). It shows the observed loss per cycle per unit length plotted (symbols) as a function of the peak magnetic field. It can be seen that these calculated values do not correspond to the measured values. Calculations were also undertaken assuming that the filaments had completely decoupled to give 75 individual conductors with cross-section of 210×20 m. It is seen that these calculations underestimate the actual measured values. It is known however, that the coupling currents which flow through the silver matrix also produce losses and therefore may be responsible for this discrepancy.